# Modified Marrone-Treanor dissociation model: formulation and verification for diatom/atom mixtures [^1] Portions of this work were presented as papers 2019-0789 and 2020-2191 at AIAA SciTech forums held in San Diego, CA. and Orlando, FL respectively

### v1

@@ 7.1 Vibrational excitation and dissociation of nitrogen and oxygen in isothermal heat baths
We verify the MMT model in scenarios where the gas initially consists only of a single diatomic species, either pure $\mathrm{N_{2}}$, or pure $\mathrm{O_{2}}$. Figure 12 shows two cases of nitrogen dissociation in isothermal, isochoric heat baths. In both these CFD-MMT calculations the gas is initially composed of pure $\mathrm{N_{2}}$ with an initial number density of $n_{0}=10^{24}\,\mathrm{m^{-3}}$ at initial vibrational temperature $T_{\mathrm{v}}\,(t\!\!=\!\!0)=300\,\mathrm{K}$, while the combined translation-rotational mode is kept artificially constant at the heat bath temperature. Macroscopic conditions imposed in the DMS reference calculations are similar to those of the CFD-MMT runs, with one major exception. The initial internal energies of all DMS particles representing the $\mathrm{N_{2}}$-gas are sampled from a Maxwell-Boltzmann distribution with equal rotational and vibrational temperatures, i.e. $T_{\mathrm{r,N_{2}}}\,(t\!\!=\!\!0)=T_{\mathrm{v,N_{2}}}\,(t\!\!=\!\!0)=300\,\mathrm{K}$, while their center-of-mass velocities are sampled from a Maxwellian distribution at the reservoir temperature $T$. In order to maintain isothermal conditions throughout the remainder of the DMS calculation, center-of-mass velocities of the $\mathrm{N_{2}}$ molecules, as well as of $\mathrm{N}$-atoms produced by dissociation, are re-sampled from their respective Maxwellians at reservoir temperature $T$ after every time step. This means that in the DMS calculations both $T_{\mathrm{r,N_{2}}}$ and $T_{\mathrm{v,N_{2}}}$ will begin to rise from their initial values and gradually approach the reservoir temperature. By contrast, in the CFD-MMT calculations it is implicitly assumed that the rotational mode starts out fully excited at the reservoir temperature and only the vibrational mode is initialized out of equilibrium.
[…]
@@
This phenomenon is the result of a temporary balance being established between two distinct processes acting simultaneously in opposite directions. On the one hand, inelastic collisions responsible for vibrational excitation of the gas drive $e_{\mathrm{N_{2}}}^{\mathrm{v}}$ upward, to its equilibrium value $e_{\mathrm{N_{2}}}^{\mathrm{v,\star}}(T)$. On the other, with every dissociation reaction a certain amount of vibrational energy is removed from the gas, causing $e_{\mathrm{N_{2}}}^{\mathrm{v}}$ to decrease. As discussed for instance in Refs. [50, 27], in the DMS solutions this quasi-steady-state dissociation phase is naturally captured by the method itself and vertical dashed lines in Figs. 12(a) and 12(b) mark its approximate beginning. The $T-T_{\mathrm{v}}$ gap in DMS occurs as a macroscopic manifestation of a temporary balance at the microscopic level between vibrational energy gain due to inelastic trajectories and nearly equal energy loss due to reactive collisions respectively. In the CFD-MMT solutions the same process is captured macroscopically by the difference between the two vibrational energy source terms of Eq. (6). Given that $T_{\mathrm{v}}<T$ throughout the entire simulated time, the vibrational relaxation term $Q_{\mathrm{v}}^{\mathrm{relax}}$ remains positive and is greatest in magnitude early on. During the vibrational excitation phase this is the dominant source term, only being counteracted by the vibrational energy-chemistry coupling term $Q_{\mathrm{chem}}^{\mathrm{v}}$ once nitrogen dissociation begins to gain prominence. After the dissociation rate has fully ramped up, $Q_{\mathrm{v}}^{\mathrm{relax}}$ and $Q_{\mathrm{chem}}^{\mathrm{v}}$ become almost equal in magnitude, but opposite in sign. Thus, according to Eq. (6) vibrational energy density, and with it $T_{\mathrm{v}}$, in the heat bath mass do not change appreciably during the QSS phase.
[…]
@@
The corresponding CFD-MMT and DMS profiles for mixture composition are shown in the lower halves of Figs. 12(a) and 12(b). At the lower-temperature condition the $\mathrm{N_{2}}$ and $\mathrm{N}$ mole fraction profiles of the CFD-MMT solution (red and orange dash-dotted lines) lead their DMS counterparts (solid lines) by a small amount. In the higher-temperature heat bath this relationship is inverted and the CFD-MMT solution lags behind its DMS counterpart by a small, equally noticeable amount. Such isothermal heat bath simulations involve extreme conditions as energy is continually being added to the system, and exact agreement between a two-temperature CFD model and DMS is not expected.
[…]
@@
Figure 12: Verification of CFD-MMT model against DMS: nitrogen dissociation in isothermal heat bath
[…]
@@
In contrast, in the 10 $10\,000\,\mathrm{K}$-oxygen heat bath, DMS-derived $\mathrm{O_{2}}$ rotational and vibrational temperatures rise over similar time scales (solid blue and red lines in Fig. 13(b)). At this condition the rotational temperature of $\mathrm{O_{2}}$ levels off at its own QSS plateau and does not reach the heat bath temperature before the simulation ends. Thus, as was seen for the 20 $20\,000\,\mathrm{K}$-nitrogen case in Fig. 12(b), in the 10 $10\,000\,\mathrm{K}$-oxygen DMS calculation a certain degree of translation-rotational nonequilibrium persists for the entire simulated time. Figure 13(b) also shows that the CFD-derived $T_{\mathrm{v}}$ profile (dash-dotted red) agrees well with the DMS-derived one (solid red), throughout the excitation and QSS phases. Both vibrational temperature profiles level off at nearly equal QSS plateaus.
[…]
@@
Figure 13: Verification of CFD-MMT model against DMS: oxygen dissociation in isothermal heat bath
@@ 7.2 Vibrational excitation and dissociation of nitrogen and oxygen in adiabatic heat baths
After the isothermal heat bath comparisons of Sec. 7.1, we perform a similar analysis for adiabatic conditions. In this scenario the heat bath’s total energy remains constant and translational temperature freely adjusts in response to internal energy redistribution processes and chemical reactions, which gradually drive the gas closer to equilibrium. These conditions allow us to better assess the MMT model’s behavior over a wide temperature range within a single calculation and it is more representative of post-shock conditions.
Four separate cases, two for nitrogen and two for oxygen are simulated using the MMT model and compared to corresponding DMS reference solutions. In the following discussion each case is identified by the heat bath’s specific enthalpy (constant throughout the simulation). The DMS results used for reference were taken from previous work (nitrogen cases 3 and 4 in Ref. [51] and “low-enthalpy” and “high-enthalpy” oxygen cases in Ref. [58]). All initial conditions are summarized in Table 4 and, as was the case for the isothermal heat baths of Sec. 7.1, they differ slightly between CFD and the equivalent DMS calculations. In the DMS calculations the translational temperature $T_{\mathrm{t}}^{\mathrm{DMS}}(t=0)$ is initialized at a higher value than the common rotational and vibrational temperatures $T_{\mathrm{r}}^{\mathrm{DMS}}(t=0)=T_{\mathrm{v}}^{\mathrm{DMS}}(t=0)$, causing the gas to undergo rapid early translation-rotational relaxation, in parallel with a slower vibrational excitation phase. By construction, in the two-temperature CFD model translational and rotational modes remain in equilibrium at all times. Thus, we must initialize the combined translation-rotational temperature to the weighted average $T_{\mathrm{t-r}}^{\mathrm{CFD}}(t=0)=\tfrac{3}{5}T_{\mathrm{t}}^{\mathrm{DMS}}(t=0)+\tfrac{2}{5}T_{\mathrm{t}}^{\mathrm{DMS}}(t=0)$ in order to match the enthalpy of the DMS case. Note that the weights $c_{v}^{\mathrm{t}}/(c_{v}^{\mathrm{t}}+c_{v}^{\mathrm{r}})=3/5$ and $c_{v}^{\mathrm{r}}/(c_{v}^{\mathrm{t}}+c_{v}^{\mathrm{r}})=2/5$ correspond to the relative energy contributions to heat capacity at constant volume of the translational and rotational modes in a fully excited diatomic gas. In each of the four cases the initial translation-rotational temperature of the CFD calculations will be lower than the DMS’s initial translational one, because the two modes in CFD begin “pre-equilibrated”. As seen in Table 4, this also has repercussions for the initial pressures in the CFD and DMS calculations.
@@
$h$ $\rho\times 10^{3}$ moles $T_{\mathrm{t}}$ $T_{\mathrm{r}}=T_{\mathrm{v}}$ $p$ $T_{\mathrm{t-r}}$ $T_{\mathrm{v}}$ $p$
[…]
nitrogen $12.3$ $2.11$ $100\%\,\mathrm{N_{2}}$ 24 0002400024\,00024 000 $300$ $15.0$ 14 5201452014\,52014 520 $300$ $9.08$
nitrogen $20.1$ $2.65$ $100\%\,\mathrm{N_{2}}$ 42 0004200042\,00042 000 $300$ $33.0$ 25 3202532025\,32025 320 $300$ $25.3$
oxygen $4.67$ $34.5$ $100\%\,\mathrm{O_{2}}$ 10 3001030010\,30010 300 $298$ $92.0$ 6 28062806\,2806 280 $298$ $56.3$
oxygen $9.81$ $20.5$ $100\%\,\mathrm{O_{2}}$ 22 5002250022\,50022 500 $298$ $120$ 13 6001360013\,60013 600 $298$ $72.7$
Figure 14 summarizes the comparisons for both nitrogen cases, the lower-enthalpy case in Fig. 14(a) and the higher-enthalpy one in Fig. 14(b). As in each of the previous figures, the upper half shows temperatures and the lower half the corresponding mole fraction profiles. All labeling conventions, line patterns and colors are carried over from Fig. 12, with two minor changes. Instead of a linear time axis, Figs. 14(a) and 14(b) both use a logarithmic scale to better highlight the early rotational and vibrational relaxation phases in the DMS calculations, while also allowing one to observe the slow decrease in temperature due to $\mathrm{N_{2}}$-dissociation later on. In addition to vibrational temperature profiles (red lines) from both methods, separate translational (solid black) and rotational (solid blue) temperatures are plotted for the DMS solution. A dash-dotted gray curve represents the CFD-derived $T_{\mathrm{t-r}}$ profile and an equivalent one for DMS is shown as a solid gray line. The latter is calculated, as just discussed, by weighing the DMS translational and rotational mode contributions according to $T_{\mathrm{t-r}}^{\mathrm{DMS}}=\tfrac{3}{5}\,T_{\mathrm{t}}^{\mathrm{DMS}}+\tfrac{2}{5}\,T_{\mathrm{r,N_{2}}}^{\mathrm{DMS}}$. During early translation-rotational relaxation captured by the DMS calculations this profile represents a fictitious average temperature, which nevertheless makes direct comparison between CFD and DMS results more straightforward.
Focusing on the lower-enthalpy case first, the CFD-derived $T_{\mathrm{t-r}}$ profile agrees remarkably well with the corresponding DMS one. It is evident that the CFD-MMT model closely matches the vibrational relaxation behavior of the DMS reference calculations. With a specific enthalpy of $12.3\,\mathrm{MJ/kg}$ this case represents the milder of the two conditions, where dissociation of nitrogen is less intense and the time scales for vibrational relaxation and chemistry are disparate enough for the gas to be nearly fully vibrationally excited, before significant dissociation begins to occur.
@@
(a) Adiabatic heat bath at $h=12.3\,\mathrm{MJ/kg}$
(b) Adiabatic heat bath at $h=20.1\,\mathrm{MJ/kg}$
Figure 14: Verification of CFD-MMT model against DMS: nitrogen dissociation in adiabatic heat bath
The two adiabatic reservoir cases with pure oxygen are shown in Fig. 15. Much like for the preceding adiabatic nitrogen cases, at the lower-enthalpy condition shown in Fig. 15(a) excellent agreement between CFD and DMS profiles is observed. More significant discrepancies between CFD and DMS solutions can be observed for the higher-enthalpy condition in Fig. 15(b). Here, a time lag between CFD and DMS for $T_{\mathrm{t-r}}$ and the species mole fraction profiles, qualitatively similar to the high-enthalpy nitrogen case of Fig. 14(b), can be observed during the first $0.7$ microseconds. The CFD solution (dash-dotted lines) slightly trails behind the DMS one (solid lines) early on, but eventually catches up. Past the first microsecond both methods are in excellent agreement up until the point where the DMS solution cuts off.
(a) Adiabatic heat bath at $h=4.67\,\mathrm{MJ/kg}$
(b) Adiabatic heat bath at $h=9.81\,\mathrm{MJ/kg}$
Figure 15: Verification of CFD-MMT model against DMS: oxygen dissociation in adiabatic heat bath

### v2

@@ 7.1 Vibrational excitation and dissociation of nitrogen and oxygen in isothermal heat baths
We benchmark the MMT model in scenarios where the gas initially consists only of a single diatomic species, either pure $\mathrm{N_{2}}$, or pure $\mathrm{O_{2}}$. Figure 14 shows two cases of nitrogen dissociation in isothermal, isochoric heat baths. In both these CFD-MMT calculations the gas is initially composed of pure $\mathrm{N_{2}}$ with an initial number density of $n_{0}=10^{24}\,\mathrm{m^{-3}}$ at initial vibrational temperature $T_{\mathrm{v}}\,(t\!\!=\!\!0)=300\,\mathrm{K}$, while the combined translation-rotational mode is kept artificially constant at the heat bath temperature. Macroscopic conditions imposed in the DMS reference calculations are similar to those of the CFD-MMT runs, with one major exception. The initial internal energies of all DMS particles representing the $\mathrm{N_{2}}$-gas are sampled from a Maxwell-Boltzmann distribution with equal rotational and vibrational temperatures, i.e. $T_{\mathrm{r,N_{2}}}\,(t\!\!=\!\!0)=T_{\mathrm{v,N_{2}}}\,(t\!\!=\!\!0)=300\,\mathrm{K}$, while their center-of-mass velocities are sampled from a Maxwellian distribution at the reservoir temperature $T$. In order to maintain isothermal conditions throughout the remainder of the DMS calculation, center-of-mass velocities of the $\mathrm{N_{2}}$ molecules, as well as of $\mathrm{N}$-atoms produced by dissociation, are re-sampled from their respective Maxwellians at reservoir temperature $T$ after every time step. This means that in the DMS calculations both $T_{\mathrm{r,N_{2}}}$ and $T_{\mathrm{v,N_{2}}}$ will begin to rise from their initial values and gradually approach the reservoir temperature. By contrast, in the CFD-MMT calculations it is implicitly assumed that the rotational mode starts out fully excited at the reservoir temperature and only the vibrational mode is initialized out of equilibrium.
In Figs. 14(a) and 14(b) we compare CFD-MMT (dotted lines) and DMS (continuous lines with symbols) results at reservoir temperatures $T=10\,000\,\mathrm{K}$ and $T=20\,000\,\mathrm{K}$ respectively. At both conditions, the CFD-MMT and DMS-derived $T_{\mathrm{v}}$ profiles rise together at almost equal rates from the initial value at $300\,\mathrm{K}$. In the $10\,000\,\mathrm{K}$-case this phase is seen to last about $2\,\mathrm{\mu s}$, while in the $20\,000\,\mathrm{K}$-case it is roughly $0.04\,\mathrm{\mu s}$ long. The vibrational temperatures then level off at near-constant values somewhat below the respective heat bath temperatures. For the first case this $T_{\mathrm{v}}$-plateau lies only about $150\,\mathrm{K}$ below $T=10\,000\,\mathrm{K}$, whereas for the second case the gap between $T=20\,000\,\mathrm{K}$ and $T_{\mathrm{v}}$ is closer to $2600\,\mathrm{K}$, about ten times greater in relative terms. The observed gaps imply that during the simulated time neither of the two cases manages to attain thermal equilibrium. Instead, after the rapid early rise in $T_{\mathrm{v}}$, the mixture settles into a QSS dissociation regime with slight thermal nonequilibrium for the remainder of simulated time.
@@
As mentioned in Sec. 5, this phenomenon is the result of a temporary balance established between two processes acting in opposite directions. First, vibrational excitation drives $e_{\mathrm{v},\mathrm{N_{2}}}(T_{\mathrm{v}})$ up toward matching its corresponding local equilibrium value $e_{\mathrm{v},\mathrm{N_{2}}}(T)$. Simultaneously, with every dissociation reaction a certain amount of vibrational energy is removed from the gas, pushing $e_{\mathrm{v},\mathrm{N_{2}}(T_{\mathrm{v}})}$ in the opposite direction. As discussed for instance in Refs. [54, 34], in the DMS solutions this quasi-steady-state dissociation phase is naturally captured by the method itself and vertical dashed lines in Figs. 14(a) and 14(b) mark its approximate beginning. The $T-T_{\mathrm{v}}$ gap in DMS occurs as a macroscopic manifestation of a temporary balance at the microscopic level between vibrational energy gain in inelastic collisions and nearly equal energy loss due to reactive collisions respectively. In the CFD-MMT solutions the same process is captured macroscopically by the difference between the two vibrational energy source terms of Eq. (6). Given that $T_{\mathrm{v}}<T$ throughout the entire simulated time, the vibrational relaxation term $w_{v}^{\mathrm{relax}}$ remains positive and is greatest in magnitude early on. During the vibrational excitation phase this is the dominant source term, only being counteracted by the vibrational energy-chemistry coupling term $w_{v}^{\mathrm{chem}}$ once nitrogen dissociation begins to gain prominence. After the dissociation rate has fully ramped up, $w_{v}^{\mathrm{relax}}$ and $w_{v}^{\mathrm{chem}}$ become almost equal in magnitude, but opposite in sign. Thus, according to Eq. (6) vibrational energy density, and with it $T_{\mathrm{v}}$, in the heat bath do not change appreciably during the QSS phase.
Although both methods produce largely similar vibrational excitation profiles, there remain slight differences. At both heat bath conditions the CFD-derived vibrational temperature (dotted red) profile leads its DMS-derived (continuous red) counterpart during the early excitation phase. The lag in the DMS $T_{\mathrm{v}}$-profiles can be attributed to competition between rotational and vibrational relaxation taking place simultaneously in the DMS calculations. Here part of the energy provided by the heat bath is channeled toward raising the gas’ rotational energy, delaying transfer to the vibrational mode in the process. These differences are especially pronounced at the higher-temperature heat bath condition, where rotational relaxation occurs at a time scale comparable to that of vibrational energy. Indeed, the $T_{\mathrm{r},\mathrm{N_{2}}}$ profile (continuous blue line) in Fig. 14(b) gradually flattens during the QSS dissociation phase, implying that in the DMS solution the rotational mode never fully equilibrates with the heat bath temperature. By contrast, in the CFD-MMT solutions the rotational mode starts out and remains fully excited at the heat bath temperature the entire time. At the lower-temperature condition the “interference” of rotational relaxation is less severe. At $T=10\,000\,\mathrm{K}$ the characteristic time scale for rotational relaxation is much shorter than that of the vibrational mode and, as seen by the rapid approach of the DMS-derived $T_{\mathrm{r},\mathrm{N_{2}}}$ profile to match $T$ in Fig. 14(a) (continuous blue line), is complete within the first microsecond. Thus, the implicit assumption of a fully excited rotational mode in the CFD-MMT solution is satisfied to a much greater degree in the DMS reference solution at this lower temperature.
@@
The corresponding CFD-MMT and DMS profiles for mixture composition are shown in the lower halves of Figs. 14(a) and 14(b). At the lower-temperature condition the $\mathrm{N_{2}}$ and $\mathrm{N}$ mole fraction profiles of the CFD-MMT solution (red and orange dotted lines) lead their DMS counterparts (solid lines) by a small amount. In the higher-temperature heat bath this relationship is inverted and the CFD-MMT solution lags behind its DMS counterpart by a small, equally noticeable amount. Such isothermal heat bath simulations involve extreme conditions as energy is continually being added to the system, and exact agreement between a two-temperature CFD model and DMS is not expected.
(a) $\mathrm{N_{2}/N}$ isothermal heat bath at $T=10\,000\,\mathrm{K}$
(b) $\mathrm{N_{2}/N}$ isothermal heat bath at $T=20\,000\,\mathrm{K}$
@@
Figure 14: Benchmarking of CFD-MMT model against DMS for nitrogen dissociation in isothermal heat bath
In Fig. 15 we report results for CFD/DMS comparisons in an $\mathrm{O_{2}}/\mathrm{O}$ mixture at conditions similar to the preceding ones for nitrogen. Since the dissociation threshold for oxygen is about half that for nitrogen ($T_{\mathrm{D}}=59\,300\,\mathrm{K}$ for $\mathrm{O_{2}}$ vs. $113\,200\,\mathrm{K}$ for $\mathrm{N_{2}}$), it readily dissociates at temperatures several thousand kelvin lower. Therefore, oxygen comparisons are carried out in isothermal heat baths at $T=5\,000\,\mathrm{K}$ and $10\,000\,\mathrm{K}$ respectively.
As seen in Figs. 15(a) and 15(b), temperature and mixture composition profiles for the $T=5\,000\,\mathrm{K}$-oxygen case qualitatively resemble those of the $10\,000\,\mathrm{K}$-nitrogen case, and a similar observation can be made for the $10\,000\,\mathrm{K}$-oxygen vs. $20\,000\,\mathrm{K}$-nitrogen heat baths. Focusing on the lower-temperature condition first, in Fig. 15(a) one sees a rapid rise in $\mathrm{O_{2}}$ vibrational temperature and approach to a QSS-phase plateau only a few hundred kelvin below the heat bath’s constant $T=5\,000\,\mathrm{K}$. The CFD-derived (dotted red) and DMS (solid red) $T_{\mathrm{v}}$ profiles lie almost on top of each other during vibrational excitation, confirming that the values for calculating $\tau_{\mathrm{O_{2}-O_{2}}}^{\mathrm{v}}$ and $\tau_{\mathrm{O_{2}-O}}^{\mathrm{v}}$ from Table 3 do an excellent job of reproducing the DMS behavior. Just as was seen for nitrogen in Fig. 14(a), the DMS-derived $\mathrm{O_{2}}$ rotational temperature in Fig. 15(a) (solid blue line) almost instantly jumps to match the heat bath temperature. This is a consequence of the characteristic time scale for rotational relaxation being so much shorter than that for vibration at these conditions.
@@
By contrast, in the $10\,000\,\mathrm{K}$-oxygen heat bath, DMS-derived $\mathrm{O_{2}}$ rotational and vibrational temperatures rise over similar time scales (solid blue and red lines in Fig. 15(b)). At this condition the rotational temperature of $\mathrm{O_{2}}$ levels off at its own QSS plateau and does not reach the heat bath temperature before the simulation ends. Thus, as was seen for the $20\,000\,\mathrm{K}$-nitrogen case in Fig. 14(b), in the $10\,000\,\mathrm{K}$-oxygen DMS calculation a certain degree of translation-rotational nonequilibrium persists for the entire simulated time. Figure 15(b) also shows that the CFD-derived $T_{\mathrm{v}}$ profile (dotted red) agrees well with the DMS-derived one (solid red), throughout the excitation and QSS phases. Both vibrational temperature profiles level off at nearly equal QSS plateaus.
Finally, the $\mathrm{O_{2}/O}$ mole fraction profiles in the bottom half of Figs. 15(a) and 15(b) exhibit behavior qualitatively very similar to the $\mathrm{N_{2}/N}$ equivalents in the bottom half of Figs. 14(a) and 14(b) respectively. At $T=5\,000\,\mathrm{K}$ the CFD-derived $\mathrm{O_{2}}$ and $\mathrm{O}$ mole fraction profiles (dotted blue lines in bottom half of Fig. 15(a)) lead their DMS counterparts (solid lines) by a small amount. At $T=10\,000\,\mathrm{K}$ this relationship is reversed, with the CFD-derived $\mathrm{O_{2}}$ and $\mathrm{O}$ mole fraction profiles slightly lagging behind their DMS counterparts. We consider deviations of this magnitude acceptable given how much the choice of a single, temperature-independent non-Boltzmann factor allows us to simplify the MMT model.
(a) $\mathrm{O_{2}/O}$ isothermal heat bath at $T=5\,000\,\mathrm{K}$
(b) $\mathrm{O_{2}/O}$ isothermal heat bath at $T=10\,000\,\mathrm{K}$
@@
Figure 15: Benchmarking of CFD-MMT model against DMS for oxygen dissociation in isothermal heat bath
@@ 7.2 Vibrational excitation and dissociation of nitrogen and oxygen in adiabatic heat baths
After the isothermal heat bath comparisons of Sec. 7.1, we perform a similar analysis for adiabatic conditions. In this scenario the heat bath’s total energy remains constant and translational temperature freely adjusts in response to internal energy redistribution processes and chemical reactions, which gradually drive the gas closer to equilibrium. These conditions allow us to better assess the MMT model’s behavior over a wide temperature range within a single calculation and it is arguably more representative of post-shock conditions.
Four separate cases, two for nitrogen and two for oxygen are simulated using the MMT model and compared to corresponding DMS reference solutions. In the following discussion each case is identified by the heat bath’s specific enthalpy (constant throughout the simulation). The DMS results used for reference were taken from previous work (nitrogen cases 3 and 4 in Ref. [33] and “low-enthalpy” and “high-enthalpy” oxygen cases in Ref. [67]). All initial conditions are summarized in Table 4 and, as was the case for the isothermal heat baths of Sec. 7.1, they differ slightly between CFD and the equivalent DMS calculations. In the DMS calculations the translational temperature $T_{\mathrm{t}}^{\mathrm{DMS}}(t=0)$ is initialized at a higher value than the common rotational and vibrational temperatures $T_{\mathrm{r}}^{\mathrm{DMS}}(t=0)=T_{\mathrm{v}}^{\mathrm{DMS}}(t=0)$, causing the gas to undergo rapid early translation-rotational relaxation, in parallel with a slower vibrational excitation phase. By construction, in the two-temperature CFD model translational and rotational modes remain in equilibrium at all times. Thus, we must initialize the combined translation-rotational temperature to the weighted average $T_{\mathrm{t-r}}^{\mathrm{CFD}}(t=0)=\tfrac{3}{5}T_{\mathrm{t}}^{\mathrm{DMS}}(t=0)+\tfrac{2}{5}T_{\mathrm{r,A_{2}}}^{\mathrm{DMS}}(t=0)$ in order to match the enthalpy of the DMS case. Note that the weights $c_{v}^{\mathrm{t}}/(c_{v}^{\mathrm{t}}+c_{v}^{\mathrm{r}})=3/5$ and $c_{v}^{\mathrm{r}}/(c_{v}^{\mathrm{t}}+c_{v}^{\mathrm{r}})=2/5$ correspond to the relative energy contributions to heat capacity at constant volume of the translational and rotational modes in a fully excited diatomic gas. In each of the four cases the initial translation-rotational temperature of the CFD calculations will be lower than the DMS’s initial translational one, because the two modes in CFD begin “pre-equilibrated”. As seen in Table 4, this also has repercussions for the initial pressures in the CFD and DMS calculations.
@@
$h$ $\rho\times 10^{3}$ moles $T_{\mathrm{t}}$ $T_{\mathrm{r},\mathrm{A_{2}}}=T_{\mathrm{v},\mathrm{A_{2}}}$ $p$ $T$ $T_{\mathrm{v}}$ $p$
mixture $\mathrm{[MJ/kg]}$ $\mathrm{[kg/m^{3}]}$ $\mathrm{[-]}$ $\mathrm{[K]}$ $\mathrm{[K]}$ $\mathrm{[kPa]}$ $\mathrm{[K]}$ $\mathrm{[K]}$ $\mathrm{[kPa]}$
nitrogen $12.3$ $2.11$ $100\%\,\mathrm{N_{2}}$ 24 00024\,000 $300$ $15.0$ 14 52014\,520 $300$ $9.08$
nitrogen $20.1$ $2.65$ $100\%\,\mathrm{N_{2}}$ 42 00042\,000 $300$ $33.0$ 25 32025\,320 $300$ $25.3$
oxygen $4.67$ $34.5$ $100\%\,\mathrm{O_{2}}$ 10 30010\,300 $298$ $92.0$ 6 2806\,280 $298$ $56.3$
oxygen $9.81$ $20.5$ $100\%\,\mathrm{O_{2}}$ 22 50022\,500 $298$ $120$ 13 60013\,600 $298$ $72.7$
Figure 16 summarizes the comparisons for both nitrogen cases, the lower-enthalpy case in Fig. 16(a) and the higher-enthalpy one in Fig. 16(b). As in each of the previous figures, the upper half shows temperatures and the lower half the corresponding mole fraction profiles. Most labeling conventions, line patterns and colors are carried over from Fig. 14, with two minor changes. Instead of a linear time axis, Figs. 16(a) and 16(b) both use a logarithmic scale to better highlight the early rotational and vibrational relaxation phases in the DMS calculations, while also allowing one to observe the slow decrease in temperature due to $\mathrm{N_{2}}$-dissociation later on. In addition to vibrational temperature profiles (red lines) from both methods, separate translational (solid black) and rotational (solid blue) temperatures are plotted for the DMS solution. A dotted gray curve represents the CFD-derived trans-rotational temperature $T$ and an equivalent one for $T_{\mathrm{t-r}}^{\mathrm{DMS}}$ is shown as a solid gray line. The latter is calculated, as just discussed, by weighing the DMS translational and rotational mode contributions according to $T_{\mathrm{t-r}}^{\mathrm{DMS}}=\tfrac{3}{5}\,T_{\mathrm{t}}^{\mathrm{DMS}}+\tfrac{2}{5}\,T_{\mathrm{r,N_{2}}}^{\mathrm{DMS}}$. During early translation-rotational relaxation captured by the DMS calculations this profile represents a fictitious average temperature, which nevertheless makes direct comparison between CFD and DMS results more straightforward.
Focusing on the lower-enthalpy case first, the CFD-derived trans-rotational temperature profile agrees remarkably well with the corresponding DMS one. It is evident that the CFD-MMT model closely matches the vibrational relaxation behavior of the DMS reference calculations. With a specific enthalpy of $12.3\,\mathrm{MJ/kg}$ this case represents the milder of the two conditions, where dissociation of nitrogen is less intense and the time scales for vibrational relaxation and chemistry are disparate enough for the gas to be nearly fully vibrationally excited, before significant dissociation begins to occur.
@@
(a) $\mathrm{N_{2}/N}$ adiabatic heat bath at $h=12.3\,\mathrm{MJ/kg}$
(b) $\mathrm{N_{2}/N}$ adiabatic heat bath at $h=20.1\,\mathrm{MJ/kg}$
Figure 16: Benchmarking of CFD-MMT model against DMS for nitrogen dissociation in adiabatic heat bath
The two adiabatic reservoir cases with pure oxygen are shown in Fig. 17. Much like for the preceding adiabatic nitrogen cases, at the lower-enthalpy condition shown in Fig. 17(a) excellent agreement between CFD and DMS profiles is observed. More significant discrepancies between CFD and DMS solutions can be observed for the higher-enthalpy condition in Fig. 17(b). Here a time lag between $T$ from CFD and $T_{\mathrm{t-r}}^{\mathrm{DMS}}$, as well as both methods’ species mole fraction profiles, qualitatively similar to the high-enthalpy nitrogen case of Fig. 16(b), can be observed during the first $0.7$ microseconds. The CFD solution (dotted lines) slightly trails behind the DMS one (solid lines) early on, but eventually catches up. Past the first microsecond both methods are in excellent agreement up until the point where the DMS solution cuts off.
(a) $\mathrm{O_{2}/O}$ adiabatic heat bath at $h=4.67\,\mathrm{MJ/kg}$
(b) $\mathrm{O_{2}/O}$ adiabatic heat bath at $h=9.81\,\mathrm{MJ/kg}$
Figure 17: Benchmarking of CFD-MMT model against DMS for oxygen dissociation in adiabatic heat bath